# Neuron Empirical Gradient: Connecting Neurons’ Linear Controllability and Representational Capacity

### v1

@@ 3.2 Understanding Neuron Controllability
2). Does only a small fraction of neurons exhibit controllability? Figure 3 shows the cumulative distribution of empirical gradient magnitudes for all neurons in PLMs. Rising curves are steady and do not converge until all neurons are present, indicating that a substantial number of neurons can alter the PLMs’ output. It demonstrates that no “decisive” knowledge neurons can control knowledge representation, while others have zero effect. Notably, Llama2 models exhibit smoother and more evenly distributed curves compared to BERTs, implying that neuron-level controllability is more widely dispersed in Llama2.
@@
3). Do neurons have polarity preference? Table 2 reports the ratios of positive/negative neurons. It shows that the number of positive neurons is nearly equivalent to negative neurons, indicating that PLMs show no preference for their neurons’ polarities. This finding suggests that enhancing or suppressing neurons should be guided by their gradient polarity rather than merely increasing or decreasing their activations (Dai et al., 2022; Wang et al., 2024b).
Figure 4: Means and variances of empirical gradient magnitude across Transformer layers on six models. The data is calculated from the average of 1000 prompts.
The analyses of 2),3) suggest the Dynamic Knowledge Store Hypothesis, which posits that The representation of specific knowledge by neurons is not a fixed state determined solely by their activations but rather a dynamic, balanced state that can be reconfigured through adjustments to their activations. For example, increasing the activations of both positive and negative neurons simultaneously can result in different activation patterns producing the same output probability. This hypothesis offers a fresh perspective compared to previous studies, such as knowledge neuron research (Dai et al., 2022; Geva et al., 2022; Yu and Ananiadou, 2024), which propose that knowledge is localized to a ‘small subset of neurons. It also contrasts with approaches that use the absolute values of neuron activations as indicators of knowledge representation (Voita et al., 2024; Geva et al., 2021). These differences highlight the need for more rigorously designed methodologies to avoid one-sided interpretations of LLM mechanisms, as focusing on a limited subset of neurons risks oversimplifying the complex and distributed nature of knowledge representation in PLMs.
In subsequent sections, we explore whether empirical gradients have the capacity to represent diverse language knowledge, termed “language skills.” If validated, this would connect knowledge representation to model output through neurons, allowing neuron-level model behavior adjustment.
@@ 4 Skill Neuron Probing using NeurGard
We have demonstrated that neurons could linearly influence output probability on factual probing tasks, showcasing their potential to manipulate model outputs. Building on this, we propose to investigate whether empirical gradients can effectively encode diverse language skills through the skill neuron probing (Wang et al., 2022). Skill neuron probing aims to locate neurons that encode the skill to solve language tasks. While previous studies explore the effectiveness of using neuron activations to identify skill neurons (Wang et al., 2022; Song et al., 2024), the representational capacity of empirical gradients is still underexplored.
@@ 4.1 Task Definition
We formulate the skill neuron probing task as follows. A dataset conveying specific language skills $\mathcal{D}$ consists of language sequence pairs, including knowledge inquiries $\mathcal{Q}=\{q_{1},...,q_{|\mathcal{T}|}\}$ and answer sequences $\mathcal{A}=\{a_{1},...,a_{|\mathcal{T}|}\}$, where arbitrary $a_{i}$ belongs to the answer candidate set $\mathcal{\hat{A}}_{\text{cands}}$. For example, in the sentiment classification task, $Q$ is the documents set, and $A$ is the ground-truth sentiment labels. We then build classifiers that take behaviors of arbitrary neuron subset $\mathcal{N}_{s}\subseteq\mathcal{N}$ as features to indicate the correct answer sequences $a_{i}$ for the knowledge inquiry $p_{i}$. $\mathcal{N}$ refers to all the neurons. [^10] We focus on intermediate outputs (neurons) of FF layers.
@@
$\quad\text{Acc}(f(\mathcal{N}_{s}),D)=\frac{1}{|D|}\sum_{i=1}^{|D|}\mathbbm{1}[f(\mathcal{N}_{s},p_{i})=a_{i}].$ (3)
Here, $f(\mathcal{N}_{s},p_{i})$ is the output of the classifier $F$ using the neuron subset $\mathcal{N}_{s}$ for the prompt $p_{i}$. $\mathbbm{1}[X=Y]$ is an indicator function that equals $1$ if $X$ matches $Y$, and 00 otherwise.
@@ 4.2 Evaluation Benchmark: MCEval8K
As skill neuron probing requires a fixed target token, it faces high computational costs due to the infinite possibility of answer sequences. We thus create a multi-choice language skill evaluation benchmark, MCEval8K, that forces PLMs to generate a single-token option label (A, B, etc.) named for category labels (A: positive, B: negative, etc.). MCEval8K encompasses 22 tasks across 6 distinct genres, conveying diverse language skills to evaluate the neurons’ representation capacity. Since tasks vary in different sizes, with some, such as cLang-8 (Rothe et al., 2021; Mizumoto et al., 2011), containing millions of data points, we standardize the evaluation by limiting each task to 8K queries. [^11] Only the Stereoset task has fewer than 8K queries due to the limited size of the original dataset. It minimizes unnecessary computational costs while ensuring consistency across tasks. We also ensure the number of ground-truth options per task is balanced to eliminate bias introduced by imbalanced classification. The skill genres and contained tasks are shown below (detailed in § C).
Linguistic: Part-of-Speech tagging on Universal Dependencies (POS) (Nivre et al., 2017), phrase-chunking on CoNLL-2000 (CHUNK) (Tjong Kim Sang and Buchholz, 2000), named entity recognition on CoNLL-2003 (NER) (Tjong Kim Sang and De Meulder, 2003) and grammatical error detection on cLang-8 dataset (GED) (Rothe et al., 2021; Mizumoto et al., 2011).
Content classification: Sentiment (IMDB) (Maas et al., 2011), topic classification (Agnews) (Zhang et al., 2015), and Amazon reviews with numerical labels (Amazon) (Hou et al., 2024).
Natural language inference (NLI): textual entailment (MNLI) (Williams et al., 2018), paraphrase identification (PAWS) (Zhang et al., 2019), and grounded commonsense inference (SWAG) (Zellers et al., 2018).
Factuality: Fact-checking (FEVER) (Thorne et al., 2018), factual knowledge probing (MyriadLAMA) (Zhao et al., 2024), commonsense knowledge (CSQA) (Talmor et al., 2019) and temporary facts probing (TempLAMA) (Dhingra et al., 2022).
Self-reflection: Examine PLMs’ internal status, including hallucination (HaluEval) (Li et al., 2023), toxicity (Toxic) (cjadams et al., 2017) and stereotype (Stereoset) (Nadeem et al., 2021) detections.
Multilinguality: We select tasks containing queries in different languages, including language identification (LTI) (Brown, 2014; Lovenia et al., 2024), multilingual POS-tagging on Universal Dependencies (M-POS) (Nivre et al., 2017), Amazon review classification (M-Amazon) (Keung et al., 2020), factual knowledge probing (mLAMA) (Kassner et al., 2021) and textual entailment (XNLI) (Conneau et al., 2018).
Figure 5: MCEval8K accuracies on Llama2-7B across tasks in zero-shot and few-shot settings, reported for Rand (random guess), TProb (token probability), and three proposed probers. The legend shows the average accuracy per method. See Table 8 and Table 9 for detailed accuracies values.
5 Neuron Gradient as Knowledge Feature
We train skill neuron probers based on NeurGrad’s estimated gradients to investigate whether and how empirical gradients encode language knowledge.
@@ 5.1 Gradient-based Skill Neuron Prober
To make prediction of $q_{i}$, we measure all polarities of $\mathbf{x}_{q_{i},a_{j}}^{n_{k}}$, where $a_{j}\in\hat{\mathcal{A}}_{\text{cands}},n_{j}\in\mathcal{N}_{s}^{*}$. The prediction of each $p_{i}$ is made as follows:
$f(\mathcal{N}_{s}^{*},p_{i})$ $=\underset{a_{j}\in\hat{\mathcal{A}}_{\text{cands}}}{\arg\max}\sum_{n_{k}\in\mathcal{N}_{s}^{*}}\mathbbm{1}[\mathbf{x}_{q_{i},a_{j}}^{n_{k}}=\bar{\mathbf{x}}^{n_{k}}]$ (4)
@@
Magnitude-based majority vote (Magn-prober) utilizes gradient magnitudes as features for a majority-vote classifier. During training, for a specific $p_{i}$ and $n_{k}$, we compare the gradients between $a\in\hat{\mathcal{A}}_{\text{cands}}$. Neurons that consistently exhibit the largest or smallest gradients for the ground truth $a_{i}$ compared to other candidates are used as skill indicators. We record each neuron’s preference for being either the largest or smallest. Neurons exhibiting more consistent behavior are assigned higher importance and identified as skill neurons. During inference, similar to Eq. 4, the prediction is made by selecting $a_{j}$ that satisfies the majority of $n_{k}\in\mathcal{N}_{s}^{*}$. This prober is designed to compare against the polarity-based prober, aiming to investigate the differences between using polarity and gradient magnitude as feature sources.
Random-forest classifier (Tree-prober) is finally introduced to understand the impact of considering the interdependency across skill neurons. We use the index (non-negative integers) of $a\in\hat{\mathcal{A}}_{\text{cands}}$ with the largest gradients as features of the neurons for training. The hyperparameter includes the number of trees (#n_trees) and layers (#n_layers) used in each tree. See more details in § D.2.
5.2 Experiment Setup
Dataset & Prompt settings Since our probing method restricts the output sequence length to 1, we carefully craft instructions and options for all datasets in MCEval8K through human effort. We evaluate both zero-shot and few-shot settings, ensuring in few-shot experiments that all candidate tokens appear once in the demonstrations to prevent majority label bias Zhao et al. (2021). See § F for the designed instructions for all tasks.
Prober settings During validation, we select the optimal neuron size for majority-vote probers from $2^{n}$ (0<=n<=13). For the random-forest prober, we report accuracy using scikit-learn’s default settings, where the optimal subset of features is selected automatically: 100 trees with no depth limitation. See § D.1 for detailed prober settings.
Model We perform skill neuron probing on Llama2-7B using three probers, all datasets in MCEval8K, and the full training set (6000) per task. For Llama2-70B, due to high cost, we probe one dataset per genre—NER, Agnews, PAWS, CSQA, HaluEval, and mLAMA—using 1,024 training examples and only train major-vote probers.
@@ 5.3 Result and Analysis
Empirical gradients encode language skills. Figure 5 shows accuracies for all tasks in MCEval8K using the Llama2-7B, with both zero- and few-shot settings. The results demonstrate that LM-Prob outperforms Rand, indicating that Llama2-7B is capable of understanding instructions and recalling skills from its parameters. We also confirm the effectiveness of our skill neuron probers in addressing language tasks. The Tree-prober outperforms LM-Prob by nearly 30%, and even the two simple major-vote classifiers outperform LM-Prob in both zero- and few-shot settings. The per-task classification accuracies in Figure 5 show that skill neurons effectively represent diverse language skills, achieving consistently high results across tasks. See Table 8,9 for accuracy values.
@@
Larger PLMs excel in skill recall. Table 3 compares accuracies of LM-Prob and Magn-prober across six tasks in the few-shot setting between Llama-7B and -70B. Llama-70B outperforms Llama-7B in both LM-Prob and skill neuron probing. However, the difference between LM-Prob and Magn-prober is smaller in Llama2-70B than in Llama2-7B, indicating the large model’s strong ability to recall knowledge from its parameters.
6 Properties of Skill Neurons
Neuron sizes Tasks
$2^{0}\sim 2^{3}$ Toxic, LTI, M-POS, FEVER, TempLAMA
$2^{9}\sim 2^{13}$ Agnews, MyriadLAMA, CSQA, mLAMA
Table 4: Optimal number of skill neurons in probers.
6.1 Representation & Acquisition Efficiency
Representational efficiency: By finding the optimal neuron size on the validation set, we observe that skill-neuron prober can achieve high accuracy with a few neurons. We summarize optimal neuron sizes for all tasks with Magn-prober in Table 4. Most tasks achieved optimal accuracy within 256 neurons, demonstrating the efficiency of empirical gradients in representing language skills. Notably, factuality tasks, such as MyriadLAMA, CSQA, and mLAMA, engage a larger number of neurons, suggesting that handling facts requires more diverse neurons, reflecting the complexity of factual understanding tasks.
Figure 6: Accuracies with varying training sizes.
6.2 Generality Across Diverse Contexts
Using PAWS as an example, we create 12 distinct contexts by varying the instructions, the selection of demonstrations, and the output token styles. By measuring the generality for different combinations, we observe that the generality for prompting settings with different instructions and demonstrations is very high (close to 1), while the generality largely decreases if target tokens are changed. The results indicate that skill neurons maintain strong generality across different inputs, including variations in instructions and demonstrations. However, this generality diminishes when the output tokens are changed. See § G for details of experimental settings and results, including 12 designed contexts and generality results.
6.3 Are Neurons Exclusive in Skill Representations?
We investigate whether skill neurons exclusively represent specific skills or can be substituted by different neuron sets. We thus build Magn-probers using various neuron sets. Specifically, we select 64 consecutive neurons from the ranked list, ordered by their importance as skill indicators (§ 5.1). [^12] We use 64-neurons units, which maintain high accuracies across tasks (§ A.2). With 352,256 neurons in Llama2-7B’s FF layers, this yields 5,504 accuracy values per task.
Figure 7 depicts the accuracies across six tasks, The result suggests that skill neurons are broadly distributed, with numerous neurons acting as skill indicators. Even when relying on less important neurons, the model’s representational ability only gradually declines. Moreover, using only the least important neurons (end of each line) still yields better performance than random guesses, underscoring the inclusivity of skill neurons (See § E.3 for inclusivity evaluation on all datasets).
Figure 8: Accuracies of Tree-probers with varying depths and trees. X-axis: logarithm of trees’ number; Y-axis: logarithm of tree depths; Z-axis: Accuracy.
6.4 Do skill neurons depend on each other?
The majority-vote probers assume independence between neurons, while the Tree-prober considers their interdependencies by building hierarchical classifiers, which advantage over the major-vote prober in Figure 5 suggests that language skills can be better represented when considering the inter-neurons dependency. To see how important interdependency is in representing language skills, we train Tree-probers with varying hyperparameters, including the number of trees and depths per tree. [^13] The numbers of trees and depths are set to $2^{N}$ and $2^{M}$, respectively, where $0\leq N\leq 10$, $1\leq M\leq 11$, and $N+M<12$. We report the resulting accuracies on PAWS, CSQA, and HaluEval in Figure 8. Their different shapes indicate that the interdependency levels required for different language skills are different. Some tasks (PAWS) prefer deep layers, while some (CSQA) prefer more trees, and some (HaluEval) require a balance between depths and trees. See § E.4 for more details.
7 Related Work
Mechanistic interpretability and knowledge attribution methods. Existing studies built the understanding of connections between knowledge and diverse modules in Transformers, such as attention heads (Clark et al., 2019; Olsson et al., 2022; Oymak et al., 2023), neurons in FF layers (Geva et al., 2021, 2022; Dai et al., 2022; Wang et al., 2024b), and the circuits within the models (Meng et al., 2022; Lieberum et al., 2023; Yao et al., 2024). They drive the development of knowledge attribution methods that assign importance scores to groups of features, indicating their relevance to the model output for a given input, including gradient-based method (Dai et al., 2022; Sundararajan et al., 2017), casual intervention methods that modify the internal status of models and observe the causal effect (Meng et al., 2022; Goldowsky-Dill et al., 2023) and automatic-tool-based methods relying on self-explanation with LLMs (Conmy et al., 2023; Singh et al., 2023). While these methods offer valuable insights into the interpretability of LLMs, such as neuron-ngram (Voita et al., 2024) or neuron-fact connections (Dai et al., 2022), they provide only qualitative measures of neuron importance, leaving the quantitative relationship between neurons and model output unexplored.
Skill neuron probing. Neurons in FF layers show the ability to convey specific skills so that using the neuron activations solely can tackle the language tasks, which these neurons are referred to as skill neurons (Wang et al., 2022; Song et al., 2024). Existing studies found that neurons can express semantic knowledge, solving tasks like sentiment classification (Wang et al., 2022; Song et al., 2024). Neurons are also discovered to represent more complex skills, including style transfer (Lai et al., 2024) and translation (Tan et al., 2024). Previous research viewed neuron activations as knowledge indicators, highlighting their representational ability but ignoring their limited influence on model output. In contrast, our empirical gradient findings provide a stronger basis for knowledge control.

### v2

@@ 5 Understanding Neurons’ Controllability
5.1 How Are NEGs Distributed?
@@
Do neurons have polarity preference? Table 2 reports the ratios of positive/negative neurons across 1000 prompts in the six PLMs, showing that the numbers of positive/negative neurons are nearly equivalent. It indicates that PLMs show no preference for their neurons’ polarities, suggesting that enhancing or suppressing neurons should be guided by their gradient polarity rather than merely increasing or decreasing their activations Dai et al. (2022). See detailed distribution analysis in § B.2,B.3.
Figure 4: Cumulative distribution of NEG magnitudes. (X-axis: the percentiles of NEG magnitudes; Y-axis: the cumulative contribution of neurons to the total sum).
Does only a few neurons exhibit strong gradients? Figure 4 shows the cumulative distribution of NEG for all neurons in PLMs. Rising curves are steady and do not converge until all neurons are present, indicating that most neurons can affect the model’s output probabilities.
5.2 Does Linearity Hold for Multi-neuron?
We conduct multi-neuron intervention experiments to investigate: can output shifts be predicted when modifying multiple neuron activations? We randomly sample $N$ neurons and enhance them by shifting their activations according to their polarities measured by NeurGrad, in which positive and negative activation shifts are applied to positive/negative neurons. We experiment on BERTbase and Llama2-7B using neuron sizes of $2^{N}$ ($0\leq N\leq 12$) across 1000 prompts.
Figure 5 shows the average Corr across all prompt-neuron pairs with an enhancement range of [0, 0.5] and a step size of 0.01. We observe that even Corr decreases as more neurons are involved due to neuron interactions, a strong correlation ($\geq$0.7) remains even with 4096 neurons in both PLMs. Additionally, Figure 5 (right y-axis) shows that involving more neurons can cause larger output shifts, suggesting that the output shifts caused by individual neurons can be accumulated. Note that despite the small average output shifts in Llama2 due to the small NEG magnitudes per neuron (see § B.2 for details), significant probability changes can still be observed when more than 1024 neurons are involved in Llama2-7B models with specific neuron combinations. Moreover, our analysis with larger enhancement ranges reveals consistent trends: increasing the shift range reduces Corr, making output shifts less predictable. See § D for experiments with different ranges.
Figure 5: Multi-neuron enhancement with range [0,0.5] with different number of neurons.
The analysis above demonstrates that modifying neuron activations, as guided by NeurGrad, enables partial prediction of output shifts. However, the number of modified neurons and the range of modifications require careful examination.
5.3 Local Linear Approximation Hypothesis
We present the Local Linearity Approximation Hypothesis to explain neuron linearity, based on three key observations: (1) expanding the shift range reduces Corr (Figure 2); (2) PLMs with more parameters diminish the importance of individual neurons, leading to higher Corr (Figure 2); (3) involving more neurons weakens linearity (Figure 5).
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a PLM, where $x\in\mathbb{R}^{n}$ represents neuron activations and $f(x)$ represents the output token probabilities. Within a constrained region, the influence of a single neuron $x_{i}$ can be locally approximated as: $f(x_{i}+\delta e)\approx f(x_{i})+\frac{\partial f}{\partial x_{i}}\delta,$ for small $\delta$, where $e_{i}$ is the unit basis vector. This follows from the first-order Taylor expansion, reflecting the model’s local differentiability with respect to neuron activations.
@@ 6 Skill Neuron Probing using NeurGard
While NEG and NeurGrad enable neuron-level modifications to model outputs, the variability of neurons’ NEG values across different prompts limits their ability to support modifications for specific types of knowledge, such as language skills. These skills often involve handling a range of prompts that require linguistic diversity. This section explores whether NEG can effectively capture general language skills linked to diverse prompts through skill neuron probing. Skill neuron probing seeks to identify neurons that encode the ability to solve language tasks. Note that prior work (Wang et al., 2022; Song et al., 2024) explored the effectiveness of using neuron activations for this purpose, yet we focus on neurons’ NEG.
@@ 6.1 Task Definition
Following Wang et al. (2022), we formulate the skill neuron probing task as follows. A dataset conveying specific language skills $\mathcal{D}$ consists of language sequence pairs, including knowledge inquiries $\mathcal{Q}=\{q_{1},...,q_{|\mathcal{T}|}\}$ and answer sequences $\mathcal{A}=\{a_{1},...,a_{|\mathcal{T}|}\}$, where arbitrary $a_{i}$ belongs to the answer candidate set $\mathcal{\hat{A}}_{\text{cands}}$. For example, in the sentiment classification task, $Q$ is the documents set, and $A$ is the ground-truth sentiment labels. We then build classifiers that take behaviors of arbitrary neuron subset $\mathcal{N}_{s}\subseteq\mathcal{N}$ as features to indicate the correct answer sequences $a_{i}$ for the knowledge inquiry $q_{i}$. $\mathcal{N}$ refers to all the neurons. [^9] We focus on intermediate outputs (neurons) of FF layers.
@@
$\quad\text{Acc}(f(\mathcal{N}_{s}),D)=\frac{1}{|D|}\sum_{i=1}^{|D|}\mathbbm{1}[f(\mathcal{N}_{s},q_{i})=a_{i}].$ (3)
Here, $f(\mathcal{N}_{s},q_{i})$ is the output of the classifier $F$ using the neuron subset $\mathcal{N}_{s}$ for the prompt $q_{i}$. $\mathbbm{1}[X=Y]$ is an indicator function that equals $1$ if $X$ matches $Y$, and 00 otherwise.
@@ 6.2 Evaluation Benchmark: MCEval8K
This section, we create a multi-choice knowledge evaluation benchmark-MCEval8K for skill neuron probing. As skill neuron probing requires a fixed target token, previous studies (Wang et al., 2022; Song et al., 2024) relied on multi-choice datasets that forces PLMs to generate a single-token option label (A, B, etc.) named for category labels (A: positive, B: negative, etc.). While we adopt a similar setup, earlier work focused on small PLMs with datasets limited to basic language understanding tasks, which are insufficient for evaluating LLMs.
To address this, we introduce MCEval8K, a diverse multi-choice benchmark covering 22 tasks across six genres, designed to assess a wide range of language skills, incorporating most datasets from previous studies. Since tasks vary in different sizes, with some, such as cLang-8 (Rothe et al., 2021; Mizumoto et al., 2011), containing millions of data points, we standardize the evaluation by limiting each task to 8K queries. [^10] Only the Stereoset task has fewer than 8K queries due to the limited size of the original dataset. It minimizes unnecessary computational costs while ensuring consistency across tasks. We also ensure the number of ground-truth options per task is balanced to eliminate bias introduced by imbalanced classification. The skill genres, tasks, and datasets information are shown below (detailed in § F).
Linguistic: Part-of-speech tagging (POS), phrase-chunking (CHUNK), named entity recognition (NER), and grammatical error detection (GED). Content classification: Sentiment (IMDB), topic classification (Agnews), and Amazon reviews with numerical labels (Amazon). Natural language inference (NLI): textual entailment (MNLI), paraphrase identification (PAWS), and grounded commonsense inference (SWAG). Factuality: Fact-checking (FEVER), factual knowledge probing (MyriadLAMA), commonsense knowledge (CSQA), and temporary facts probing (TempLAMA). Self-reflection: Examine PLMs’ internal status, including hallucination (HaluEval), toxicity (Toxic), and stereotype (Stereoset) detections. Multilinguality: We select tasks with multilingual queries, including language identification (LTI), multilingual POS-tagging on Universal Dependencies (M-POS), Amazon review classification (M-Amazon), factual knowledge probing (mLAMA) and textual entailment (XNLI).
Figure 6: MCEval8K accuracies on Llama2-7B across tasks in zero-shot and few-shot settings, reported for Rand (random guess), TProb (token probability), and two proposed probers. Legends show average accuracies.
7 NEG as Knowledge Feature
We train skill neuron probers based on NeurGrad’s estimated gradients to investigate whether and how NEG encodes language knowledge.
@@ 7.1 Gradient-based Skill Neuron Prober
To make prediction of $q_{i}$, we measure all polarities of $\mathbf{x}_{q_{i},a_{j}}^{n_{k}}$, where $a_{j}\in\hat{\mathcal{A}}_{\text{cands}},n_{j}\in\mathcal{N}_{s}^{*}$. The prediction of each $q_{i}$ is made as follows:
$f(\mathcal{N}_{s}^{*},q_{i})$ $=\underset{a_{j}\in\hat{\mathcal{A}}_{\text{cands}}}{\arg\max}\sum_{n_{k}\in\mathcal{N}_{s}^{*}}\mathbbm{1}[\mathbf{x}_{q_{i},a_{j}}^{n_{k}}=\bar{\mathbf{x}}^{n_{k}}]$ (4)
@@
Magnitude-based majority vote (Magn-prober) utilizes gradient magnitudes as features for a majority-vote classifier. During training, for a specific $q_{i}$ and $n_{k}$, we compare the gradients between $a\in\hat{\mathcal{A}}_{\text{cands}}$. Neurons that consistently exhibit the largest or smallest gradients for the ground truth $a_{i}$ compared to other candidates are used as skill indicators. We record each neuron’s preference for being either the largest or smallest. Neurons exhibiting more consistent behavior are assigned higher importance and identified as skill neurons. During inference, similar to Eq. 4, the prediction is made by selecting $a_{j}$ that satisfies the majority of $n_{k}\in\mathcal{N}_{s}^{*}$. This prober is designed to compare against the polarity-based prober, aiming to examine whether NEGs’ magnitude can bring additional skill information compared to polarity.
7.2 Experimental Setup
Dataset & Prompt Since our probing method restricts the output sequence length to one, we carefully craft instructions and options for all datasets in MCEval8K through human effort. We evaluate both zero-shot and few-shot settings, ensuring in few-shot experiments that all candidate tokens appear once in the demonstrations to prevent majority label bias Zhao et al. (2021). See § J for the designed instructions for all tasks.
Prober During validation, we select the optimal neuron size for majority-vote probers from $2^{n}(0\leq n\leq 13)$. See § G for detailed prober settings.
Model We perform skill neuron probing on Llama2-7B using three probers, all datasets in MCEval8K, and the full training set (6000) per task. For Llama2-70B, due to high cost, we probe one dataset per genre—NER, Agnews, PAWS, CSQA, HaluEval, and mLAMA—using 1024 training examples and only train major-vote probers.
@@ 7.3 Result and Analysis
Empirical gradients encode language skills. Figure 6 shows accuracies for all tasks in MCEval8K using the Llama2-7B, with both zero- and few-shot settings. The results demonstrate that LM-Prob outperforms Rand, indicating that Llama2-7B is capable of understanding instructions and recalling skills from its parameters. We also confirm the effectiveness of our skill neuron probers in addressing language tasks. The two simple major-vote classifiers outperform LM-Prob in both settings. The per-task classification accuracies in Figure 6 show that skill neurons effectively represent diverse language skills, achieving consistently high results across tasks. See Table 7,8 for accuracy values.
@@
Larger PLMs excel in skill recall. Table 3 compares LM-Prob and Magn-prober across six tasks in the few-shot setting between Llama-7B and -70B. Llama-70B outperforms Llama-7B in both LM-Prob and skill neuron probing. However, the difference between LM-Prob and Magn-prober is smaller in Llama2-70B than in Llama2-7B, indicating the large model’s strong ability to recall knowledge from its parameters. See § H for further analysis on the properties of skill neurons: efficiency, generality, and inclusivity.